OpenStudy (anonymous):

Hey, how would i find the cosine of the angle between the plane z−3x+2y = 11 and the y-axis? Thanks for helps =)

OpenStudy (anonymous):

@AliceMath99 Write the equation of the plane in standard from -3x+2y+z=11. The easiest way to go about this problem is to do the following: First, find a vector that parallels the y-axis. The vector u = (0,1,0) should work nicely. Next, the find a vector running through the plane -3x+2y+z=11. To do this, pick two points (x1,y1,z1) and (x2,y2,z2) that satisfy the plane equation. I chose (0,0,11) and (0,5,1). Then, a vector v that lies in the plane would be given as v = (0,0,11) - (0,5,1) = (0,-5,10) Then, remember that an alternative way to calculuate the dot product is given as u*v = ||u||*||v||*cosθ where θ is the angle between the two vectors. So, we have u*v = (0,1,0)*(0,-5,10) = 0*0+1*(-5)+0*10 = -5 We also have that ||u|| = √(02+12+02) = 1 and ||v|| = √(02+(-5)2+102) = √125 = 5√5 Hence, by the alternative formula for the dot product, we have -5 = 1*5√5 * cosθ and so cosθ = -5/(5√5) = -1/√5 or cosθ = -√(5)/5 ≈-0.4472